# FragNet: A Graph Neural Network for Molecular Property Prediction with Four Layers of Interpretability

### v1

@@ 2.2.2 Substructure Analysis
In the following sections, we perform a quantitative study of the highly impactful substructures for predicting solubility, lipophilicity, and cancer drug response. To identify these substructures, we first filter the model predictions to those with low error to focus on cases where the model is correctly inferring the structure-property relationships. We then identify substructures which appear in multiple molecules and average their attention weights across all the predictions. Substructures which have high average attention across the predictions are likely to have a significant impact on the predicted property.
@@
Figure 3 illustrates the atoms with largest average attention weights for solubility, lipophilicity, and cancer drug response (CDR) prediction for molecules with absolute errors less than 0.1. It is observed that Nitrogen (N) and Carbon (C) play dominant roles in all three scenarios, followed by O and S. Compared to solubility and lipophilicity, F and Cl have a more significant impact on cancer drug response prediction.
Highly weighted bonds in compounds making absolute prediction errors less than 0.1 are shown in Figure 4. The significance of the carbon-carbon (C-C) bond is greater for solubility and lipophilicity prediction than for cancer drug response prediction. It is unsurprising that the carbon-oxygen (C-O) bond is more critical for solubility prediction than for lipophilicity prediction, primarily because of the polarity of the C-O bond. Polar substances are more soluble in polar solvents like water.
@@
We find that the fragments containing N are among the high weight fragments ($[*]N([*])[*]$) and high contribution value fragments ($CN(C)[*]$, $N[*]$). Nitrogen containing functional groups are known to be hydrophilic. Agreeing with the known chemistry, -OH which is also a well known hydrophilic functional group that has on average a positive contribution value.
The compounds containing $[Cl-]$ are salt structures. Our analysis reveals that Cl- is either the most significant or the second most significant fragment in all but one of the compounds containing Cl- in our study set according to the attention weights. Additionally, these compounds contain N+, which ranks among the two most significant substructures in 8 out of the 14 compounds. The fragment weights for these 14 compounds are illustrated in Figure S1. Considering the chemical properties, Cl- is likely to interact predominantly with the positively charged (N) atom. Consequently, the model’s identification of these structures as the most important underscores its focus on the relevant substructures.
Despite the presence of the polar C=O group, the substructures $O=C([*])[*]$ and $O=C([*])C[*]$ also have most of their contribution values between -0.5 and 0. This is likely because of the other fragments attached to these substructures. Upon further inspection, we found that the contribution of $O=C([*])[*]$ is positive when one of the substructures attached to it is OH or NH2 (see Figure S2). This result is consistent with the well-known fact that the carboxylic acid group (O=C(C)OH) is hydrophilic. When we inspected the compounds where $O=C([*])[*]$ has negative contributions, we find that in majority of them O is attached to one end of $O=C([*])[*]$ and a five membered ring system to the other end (structures are shown in Figure S3).
@@
The highly weighted fragments for cancer drug response prediction are shown in Figure 7(a). Among them are amino groups (–NH2, –NR2), carbonyl (C=O) groups, ether oxygens, hydroxyl groups and aromatic rings, which are known to be important building blocks of drug molecules [20, 19].
Because of their polarity, amino groups increase water solubility and participate in hydrogen bonding and ionic interactions with amino acid residues in proteins, nucleic acids, and other biomolecules. Studies have shown that a compound conjugating with an amino acid results in the improvement of its pharmacological activity[21, 22]) and water solubility[23, 24], and the reduction of its cytotoxicity[25].
The presence of a carbonyl (C=O) group can improve the molecular recognition process, allowing the drug to directly fit into the active sites of the target[26]. Carbonyl groups can enhance the stability of the drug molecule due its ability to form hydrogen bonds. This network of hydrogen bonds helps to prevent premature degradation and ensure that the drug reaches its site of action intact. The chemical reactivity of carbonyl groups can play a significant role in the mechanism of drug action as well[26].
@@
Apart from contributing to improve solubility, due to the electron donating and accepting ability of ether oxygens, the presence of ether groups can also increase the bioavailability of a drug[19]. Ether oxygens (-O-) can also bolster the stability of drug molecules, reducing the chances of degradation before the drug reaches its intended site of action. For instance, in penicillin V, the ether oxygen can attract the lone pair of electrons from the oxygen of the side chain carbonyl group. This withdrawal of electrons prevents the lone pair from attacking the carbonyl carbon of the $\beta$-lactam ring[19].
The presence of hydroyl (-OH) group can increase the solubility and thereby the oral absorption of drug molecules. Hydroxyl groups also affect the metabolism, duration and the selectivity[19]. Due to their ability to engage in hydrogen bonding, hydroxyl groups can be employed to modify the sites where hydrogen bonds are formed within a drug molecule. This capability allows for the regulation of both the binding affinity and specificity of the drug towards its biological target[20].
@@
In Figure 7 (b), fragments associated with the five largest and smallest contribution values in our test dataset are shown. Notably, the hydrophobic fragments also contribute negatively to the cancer drug response prediction.
To underscore the significance of identifying useful substructures in cancer drug design, Figure 8 presents data on four drugs tested on the same BRCA[27, 28] cell line type. Among these, three drugs exhibit high response values (approximately 4.1 to 4.5), while the fourth drug shows a notably lower response value of around -0.66. Analysis reveals that the drugs with high response values have relatively smaller negative contribution values compared to the drug with the low response value. Specifically, the average contributions across fragments for the highly responsive drugs are -0.089, -0.094, and -0.095, while the low response drug has an average contribution of -0.24. Additionally, the highly responsive drugs contain fragments with substantially larger positive contributions compared to the low response drug. The positive contribution values for drugs (a), (b), and (c) are 0.40355, 0.16833, and (0.155, 0.020) respectively, whereas the low response drug shows positive contribution values of just 0.007, 0.007, and 0.001. This indicates that in computational drug design, FragNet’s contribution scoring can be utilized to identify initial fragments with positive contributions, thus aiding in the design of more effective cancer drugs.
@@
The highly weighted fragment-fragment connections for predicting solubility, lipophilicity, and drug response are shown in Figure 9. In most cases, these connections involve oxygen (O) or nitrogen (N) atoms. The presence of N-H connections is notably important for lipophilicity prediction. Additionally, C=O connections are also observed to be crucial for determining all three properties. In particular, the interaction between $[*]CC[*]$ and oxygen is significant for both solubility and drug response predictions. Non-covalent interactions between $CC[*]$ and substructures containing halogen atoms are also noteworthy, as observed in the solubility datasets. Upon close inspection, we found that in almost all cases, these ethyl groups are attached to electron-donating N or O atoms. Therefore, we believe these interactions signify non-covalent interactions between the electronegative group and the anionic substructure. Also, because of the electron donating nature of the ethyl groups, it is possible that there is an attraction between Cl- and slightly positively charged ethyl groups. This ethyl halogen interaction is also observed for when Cl- is the anion.
@@ 2.3.1 Model Embeddings
To further demonstrate the effectiveness of FragNet’s representations in distinguishing different classes of molecules, we present UMAP results in Figure 10 which show the 2D projection of the embeddings from the layer immediately preceding the final linear layer of the model. The embeddings from this layer should include a condensed representation of the molecular structure that encodes the information necessary for predicting the molecular property. Therefore, the groupings of molecules in this space provides insight into the reasoning used by the model to infer the target property. For solubility, we observe that some molecules containing hydrophilic and hydrophobic fragments are located in separate regions in the embedding space.
@@ 2.3.2 Validations using Density Functional Theory (DFT) calculations
To further investigate the interpretability of FragNet, we calculated the electrostatic surface potentials of some molecules that had the FragNet model trained on solubility predicted with high accuracy using Density Functional Theory (DFT). Regions with electrostatic surface potentials (absolute value) greater than 10 were designated as polar areas, while those with potentials less than 10 were classified as non-polar areas. For each molecule, the ratio of the non-polar area to the polar area (NPR) was then calculated for the fragments. A larger NPR indicates a fragment with lower expected polarity. Therefore, we expect fragments with high NPR values to have negative or small contribution values. We observe this trend is in some molecules, as shown in Figure 11. This observation is prominent in highly hydrophobic molecules. Specifically, in molecules where the total NPR is greater 1, the average fragment contribution is -1.19, and in molecules where the total NPR is less than 1 the average fragment contribution is -0.35. In Table S1, we provide the fragment contribution and NPR values for 16 molecules.
@@ 3 Conclusion
4 Methods
FragNet Architecture
The importance of node j and the edge connecting nodes $i$ and $j$ to node $i$ is defined by $\alpha_{i,j}^{bond-graph}$ in Equation (2).
where,
The updated node features of the bond are calculated as,
$h_{i}^{bond-graph}=\sum_{j\in N(i)}\alpha_{i,j}^{bond-graph}\textbf{W}h_{j,0}^{bond-graph}$ (4)
We then use these $h_{i}^{bond-graph}$ as the edge features of the atom graph. As in the case of bond graph, the attention coefficients of the atom graph will be computed considering both node features and edge features. The initial node features of the atom graph are based on the atomic properties like atom type, number of valence electrons, etc. For the message passing in the atom graph, we do consider self edges (the edges connecting an atom to itself). The edge properties of the self edges are initialized as a vector of zeros having the same dimension as $h_{i}^{bond-graph}$.
Our model also considers a graph of molecular fragments. We use BRICS fragmentation to identify the fragments. In the case of regular molecules (not salts or complexes), the edge between two molecular fragments is a regular covalent bond. In the case of molecules with isolated fragments (like salts and complexes), where regular covalent bonds do not exist between some fragments, we create virtual bonds connecting each such isolated fragment to every other fragment (illustrated in Figure 1 [c] ). There are some molecules which cannot be fragmented. For such molecules we create a self-connection connecting itself to itself to facilitate message passing.
The initial fragment features are calculated by summing the features of the atoms belonging to the fragments.
We then use graph attention mechanism to update fragment features $h_{i,0}^{fragment-graph}$ similar to how we updated atom features of the atom graph. In order to achieve this task, we also have to consider a graph of connections, similar to the bond graph. The node features of the connection graph are one-hot encoded connection types. This connection type could be single, double, triple, aromatic, ‘self connection’, or ‘virtual connection’. The edge attributes of the connection graph are obtained by summing the node features of the nodes at the either side of the edge. The graph attention coefficients for the connection graph are computed using a expression similar to Equation 2. Then these coefficients are used to update the node features of the connection graph, which are the connections of the fragment graph. The updated node features of the fragment connection graph are used to initialize the edge features of the fragment graph.
Features
The atomic and bond features were created using the methods implemented in the RDKit package[29].
As atomic features we consider Atomic Number, Implicit Valence, Formal Charge, Number of Radical Electrons’, Hybridization, Is the atom Aromatic, Is the atom in a Ring, Total Number of H atoms, Chirality.
As the bond features we consider the Bond Type (SINGLE, DOUBLE, TRIPLE, AROMATIC), Is the bond Conjugated, Is the bond in a Ring, Stereo nature (STEREOANY, STEREOZ, STEREOE, STEREONONE), and Bond Direction (BEGINWEDGE, BEGINDASH, ENDDOWNRIGHT,ENDUPRIGHT, NONE).
Data
We leverage several datasets from MoleculeNet[18] to evaluate the performance of the FragNet model. The datasets include ESOL[5], Lipophilicity, Clintox[30, 31, 32], SIDER[33, 34] and Tox21. CEP dataset was obtained following the instructions given by [35]
For substructure studies, we used the PNNL solubliity dataset, the Lipophilicity dataset, and the CDRP dataset. The Lipophilicity dataset is the same that was used for model performance evaluation downloaded from MoleculeNet. The PNNL solubility dataset is a combination of [36], Cui[37] and SOMAS[38] datasets. The CDRP dataset was obtained from the work of [39].
Pretraining and Finetuning
We pretrained FragNet using 2,054,100 molecules from the dataset used by Uni-Mol[40]. The pretraining is based on the method outlined in the reference [41] and focuses on minimizing losses based on the predictions of bond length, bond angle, dihedral angle, and conformational energy of the molecule.
Hyperparameter tuning was performed using the Optuna package [42]. For small datasets we used a regression head with four layers. The number of nuerons in each layer were allowed to take values from 64 to 2048 in steps of 64. We also tuned the activation function (with options relu, silu, gelu, celu, selu, rrelu, relu6, prelu, leakyrelu) and batch size (with options 16, 32, 64, 128).
DFT calculation
Density functional theory (DFT) calculations were performed using the B3LYP[43, 44] functional with ORCA package.[45] The geometries of the molecules were optimized with def2-SVP basis set. Single point energies were calculated with def2-TZVP basis set. Vibrational frequencies were calculated for validation of stable configuration. An effect of implicit solvent model with water was included via Conductor-like Polarizable Continuum Model (CPCM) model.[46] The surface charge potentials were calculated by Multiwfn software.[47]
@@ 4 Methods
The code is available at https://github.com/pnnl/FragNet

### v2

@@ 3.2.2 Substructure Analysis
In the following sections, we perform a quantitative study of the highly impactful substructures for predicting solubility, lipophilicity, and cancer drug response. To identify these substructures, we first filter the model predictions to those with low error to focus on cases where the model is correctly inferring the structure-property relationships. This filtration is important to ensure that the attention weights are reliable in interpreting the model predictions. We then identify substructures which appear in multiple molecules and average their attention weights across all the predictions. Substructures which have high average attention across the predictions are likely to have a significant impact on the predicted property.
@@
In Figure 4 we illustrate the atoms with the largest average attention weights for solubility, lipophilicity, and cancer drug response (CDR) prediction in molecules with absolute errors less than 0.1. In examining the learned atom-based importance weights for the three predictive tasks, several salient trends emerge that correspond well with established principles in medicinal and physical chemistry 37. Heteroatoms—particularly nitrogen, oxygen, and sulfur—consistently receive higher weights across all three models, in line with their known roles in modulating hydrogen bonding, polarity, and reactivity 38. These properties strongly influence molecular solubility and pharmacologically relevant features such as absorption, distribution, metabolism, and excretion (ADME). Carbon, unsurprisingly, also shows high importance in all tasks due to its role as the primary structural scaffold in organic compounds, thereby dictating the molecular framework and overall hydrophobic/hydrophilic balance 39. Another prominent observation is the relatively higher weighting of chlorine and fluorine for both lipophilicity and cancer-drug response predictions, consistent with longstanding medicinal chemistry strategies that use halogenation to fine-tune lipophilicity, metabolic stability, and target engagement 40. Finally, hydrogen appears less influential for aqueous solubility, likely because it is overshadowed by the effect of more polar heteroatoms; yet it remains non-trivial for lipophilicity and biological response, underscoring how subtle alterations in hydrogen content can modulate binding interactions and overall drug-like behavior 38.
In Figure 5, we show the highly weighted bonds for compounds whose predictions have absolute errors less than 0.1. We notice that C–C and C–N bonds receive consistently high weights for predicting solubility, lipophilicity , and cancer drug response. This trend is unsurprising, as most drug-like small molecules contain a primarily carbon-based skeleton alongside strategically placed nitrogen atoms, which can strongly influence pharmacokinetic and pharmacodynamic profiles by modulating hydrogen-bonding capacity, basicity, and metabolic stability37. Notably, bonds that include other heteroatoms, such as C–O or C–S, also appear among the top weights in all three models, underscoring the well-recognized influence of polar functional groups on aqueous solubility 41. The presence of bonds to halogens (e.g., C–F or C–Cl) shows moderate but still significant influence on cancer drug response prediction, reflecting the common medicinal-chemistry practice of introducing halogens to tune lipophilicity, improve metabolic stability, and enhance binding affinity 42.
@@
We find that the fragments containing N are among the high weight fragments ($[*]N([*])[*]$) and high contribution value fragments ($CN(C)[*]$, $N[*]$). Nitrogen containing functional groups are known to be hydrophilic. Agreeing with the known chemistry, -OH which is also a well known hydrophilic functional group has on average a positive contribution value.
The compounds containing $[Cl^{-}]$ are salt structures. Our analysis reveals that $Cl^{-}$ is either the most significant or the second most significant fragment in all but one of the compounds containing $Cl^{-}$ in our study set according to the attention weights. Additionally, these compounds contain N+, which ranks among the two most significant substructures in 8 out of the 14 compounds. The fragment weights for these 14 compounds are illustrated in Figure S1. Considering the chemical properties, $Cl^{-}$ is likely to interact predominantly with the positively charged (N) atom. Consequently, the model’s identification of these structures as the most important underscores its focus on the relevant substructures.
Despite the presence of the polar C=O group, the substructures $O=C([*])[*]$ and $O=C([*])C[*]$ also have most of their contribution values between -0.5 and 0. This is likely because of the other fragments attached to these substructures. Upon further inspection, we found that the contribution of $O=C([*])[*]$ is positive when one of the substructures attached to it is OH or NH2 (see Figure S2). This result is consistent with the well-known fact that the carboxylic acid group (O=C(C)OH) is hydrophilic. When we inspected the compounds where $O=C([*])[*]$ has negative contributions, we find that in majority of them, one end of $O=C([*])[*]$ is attached to a ring system (structures are shown in Figure S2).
@@
The highly weighted fragments for cancer drug response prediction are shown in Figure 8(a). Among them are amino groups (–NH2, –NR2), carbonyl (C=O) groups, ether oxygens, hydroxyl groups and aromatic rings, which are known to be important building blocks of drug molecules 43, 36.
Because of their polarity, amino groups increase water solubility and participate in hydrogen bonding and ionic interactions with amino acid residues in proteins, nucleic acids, and other biomolecules. Studies have shown that a compound conjugating with an amino acid results in the improvement of its pharmacological activity44, 45 and water solubility46, 47, and the reduction of its cytotoxicity48.
The presence of a carbonyl (C=O) group can improve the molecular recognition process, allowing the drug to directly fit into the active sites of the target49. Carbonyl groups can enhance the stability of the drug molecule due its ability to form hydrogen bonds. This network of hydrogen bonds helps to prevent premature degradation and ensure that the drug reaches its site of action intact. The chemical reactivity of carbonyl groups can play a significant role in the mechanism of drug action as well49.
@@
Apart from contributing to improve solubility, due to the electron donating and accepting ability of ether oxygens, the presence of ether groups can also increase the bioavailability of a drug36. Ether oxygens (-O-) can also bolster the stability of drug molecules, reducing the chances of degradation before the drug reaches its intended site of action. For instance, in penicillin V, the ether oxygen can attract the lone pair of electrons from the oxygen of the side chain carbonyl group. This withdrawal of electrons prevents the lone pair from attacking the carbonyl carbon of the $\beta$-lactam ring36.
The presence of hydroyl (-OH) group can increase the solubility and thereby the oral absorption of drug molecules. Hydroxyl groups also affect the metabolism, duration and the selectivity36. Due to their ability to engage in hydrogen bonding, hydroxyl groups can be employed to modify the sites where hydrogen bonds are formed within a drug molecule. This capability allows for the regulation of both the binding affinity and specificity of the drug towards its biological target43.
@@
In Figure 8 (b), fragments associated with the five largest and smallest contribution values in our test dataset are shown. Notably, the hydrophobic fragments contribute negatively to the cancer drug response prediction.
To underscore the significance of identifying useful substructures in cancer drug design, Figure 9 presents data on four drugs tested on the same BRCA50, 51 cell line type. Among these, three drugs exhibit high response values (approximately 4.1 to 4.5), while the fourth drug shows a notably lower response value of around -0.66. Analysis reveals that the drugs with high response values have relatively smaller negative contribution values compared to the drug with the low response value. Specifically, the average contributions across fragments for the highly responsive drugs are -0.089, -0.094, and -0.095, while the low response drug has an average contribution of -0.24. Additionally, the highly responsive drugs contain fragments with substantially larger positive contributions compared to the low response drug. The positive contribution values for drugs (a), (b), and (c) are 0.40355, 0.16833, and (0.155, 0.020) respectively, whereas the low response drug shows positive contribution values of just 0.007, 0.007, and 0.001. This indicates that in computational drug design, FragNet’s contribution scoring can be utilized to identify initial fragments with positive contributions, thus aiding in the design of more effective cancer drugs.
@@
The highly weighted fragment-fragment connections for predicting solubility, lipophilicity, and drug response are shown in Figure 10. In most cases, these connections are from oxygen or nitrogen to a carbon-based substituent group. This is unsurprising, given that both oxygen and nitrogen are significantly more electronegative than carbon, producing polar bonds that enable dipole–dipole interactions and, often, hydrogen bonding with polar solvents.
The presence of connections with N-H is notably important for lipophilicity prediction. The covalent bond between a nitrogen atom in an N-H group and a substituent group can significantly alter the electronic environment of the molecule. If the substituent group is electron-withdrawing or electron-donating, it changes the electron density on the nitrogen and the overall polarity of the molecule. This, in turn, affects how the molecule interacts with polar versus nonpolar environments, directly impacting lipophilicity. The substituent group can also enable or prevent intramolecular hydrogen bonding or other non-covalent interactions, which can “hide” polar groups from the solvent, effectively increasing lipophilicity by reducing exposed polarity.
Additionally, connections with C=O are also observed to be important for determining all three properties. In particular, the interaction between $[*]CC[*]$ and oxygen is significant for both solubility and drug response predictions. Non-covalent interactions between $CC[*]$ and substructures containing halogen atoms are also noteworthy, as observed in the solubility datasets. Upon close inspection, we found that in almost all cases, these ethyl groups are attached to electron-donating N or O atoms. Therefore, we believe these interactions signify non-covalent interactions between the electronegative group and the anionic substructure. Also, because of the electron donating nature of the ethyl groups, it is possible that there is an attraction between Cl- and slightly positively charged ethyl groups. This ethyl halogen interaction is also observed for when Cl- is the anion.
@@ 3.3 Additional Analysis
3.3.1 Fragments in inaccurately predicted molecules
To identify molecular fragments that the model finds challenging to predict accurately, we computed the fragment weights and contributions for the top 20% of molecules exhibiting the largest prediction errors. Subsequently, we categorized these molecules based on their constituent fragments and calculated the average prediction errors, fragment weights, and fragment contributions for each fragment group.
Fragment Error Weight Contr. In low err. In high err.
$[*]CC[*]$ 3.00 4.38 -0.01 12.83 7.76
$[*]N[*]$ 2.17 4.68 -0.01 13.58 15.53
$O[*]$ 2.13 1.61 0.06 31.32 30.12
$[*]N([*])[*]$ 2.11 7.21 0.00 6.42 5.28
$[*]C[*]$ 1.97 4.38 -0.06 3.40 8.07
$Cl$ 1.93 6.10 0.11 8.68 4.66
$CO[*]$ 1.78 0.93 -0.13 9.06 9.63
$[*]O[*]$ 1.78 4.37 -0.04 28.68 20.81
$[Na+]$ 1.76 4.21 0.25 0.00 5.90
$N[*]$ 1.69 1.94 0.13 8.30 11.18
Table 3: Fragments Identified in high-error compounds. ‘Error’ represents the average absolute error; ‘Weight’ denotes the average fragment attention weight; ‘Contr.’ indicates the average fragment contribution to solubility. ‘In low err.’ shows the percentage of low-error molecules containing the fragment, while ‘In high err.’ denotes the percentage of high-error molecules containing the fragment.
In Table 3, we present the fragments found in molecules with high prediction errors, ordered by the magnitude of the errors associated with the molecules they belong to. Specifically, we display the fragments corresponding to the ten largest average errors. To further refine our analysis, we consider fragments with average weight values equal to or above 4. The fragments that meet this criterion include $[*]CC[*]$, $[*]N[*]$, $[*]N([*])[*]$, $[*]C[*]$, $Cl$, $[*]O[*]$, and $[Na+]$.
It is noteworthy that some of these fragments are also present in molecules that were predicted with high accuracy. To identify any fragments predominantly found in high-error molecules, we calculated the percentage of molecules containing each fragment in both low-error and high-error groups. Here, low-error molecules refer to those discussed in Section 3.2.2 with prediction errors of less than 0.1, while high-error molecules are the top 20% of molecules with the largest prediction errors considered in this section. These findings are depicted in the right panel of Table 3.
From the analysis, we observe that $[Na+]$ is a fragment exclusively found in high-error molecules. Consequently, compounds containing $[Na+]$ can be regarded as challenging for the model to predict accurately. This insight suggests that further analysis is warranted to identify the underlying reasons and potential improvements needed to enhance the model’s predictive performance for such molecules.
@@ 3.3.2 Model Embeddings
To further demonstrate the effectiveness of FragNet’s representations in distinguishing different classes of molecules, we present UMAP results in Figure 11 which show the 2D projection of the embeddings from the layer immediately preceding the final linear layer of the model. The embeddings from this layer should include a condensed representation of the molecular structure that encodes the information necessary for predicting the molecular property.
We observe that there are distinct regions in the embedding space where molecules with specific fragments are densely populated. This indicates that the learned embeddings for these molecules are significantly influenced by these substructures. Hence, the groupings of molecules in this space provides insight into the reasoning used by the model to infer the target property. For example, with regard to solubility prediction, some molecules containing hydrophilic and hydrophobic fragments are located in separate regions in the embedding space. It should also be noted that the properties of these molecules are not solely determined by a specific functional group attached to them. This is the reason why we notice that molecules containing different functional groups are not in completely isolated clusters.
@@ 3.3.3 Validations using Density Functional Theory (DFT) calculations
To further investigate the interpretability of FragNet, we calculated the electrostatic surface potentials of some molecules that had the FragNet model trained on solubility predicted with high accuracy using Density Functional Theory (DFT). The details on the DFT calculations are provided in the Supporting Information section. Regions with electrostatic surface potentials (absolute value) greater than 10 were designated as polar areas, while those with potentials less than 10 were classified as non-polar areas. For each molecule, the ratio of the non-polar area to the polar area (NPR) was then calculated for the fragments. A larger NPR indicates a fragment with lower expected polarity. Therefore, we expect fragments with high NPR values to have negative or small contribution values. We observe this trend is in some molecules, as shown in Figure 12. This observation is prominent in highly hydrophobic molecules. Specifically, in molecules where the total NPR is greater 1, the average fragment contribution is -1.19, and in molecules where the total NPR is less than 1 the average fragment contribution is -0.35. In Table S1, we provide the fragment contribution and NPR values for 16 molecules.
@@ 3.3.4 Interactive Browser Application
4 Future Work
The recent integration of large language models (LLMs) in explainable AI represents a significant leap towards generating more natural and accessible explanations52, 53. LLMs, combined with techniques like Retrieval Augmented Generation (RAG) and tool use, leverage vast repositories of knowledge to provide contextually rich and coherent explanations We envision future developments where FragNet’s interpretability mechanisms are enhanced by integrating LLMs and Vision Language Models to generate explanations that are both scientifically rigorous and easily comprehensible. Furthermore, FragNet can be combined with LLMs for molecular property optimization workflows. In Figure 14, we demonstrate how FragNet-predicted contribution values can be utilized to prompt an LLM to design molecules with improved property values.
Figure 14: Prompting Claude 3.5 Sonnet with and without FragNet’s contribution data to create a molecule with improved solubility. The new fragments in the LLM-modified molecule are highlighted in green.
@@ 5 Conclusion
[…]
@@
G.P. conceived the idea, led the project, developed the model, analyzed the results and wrote the manuscript. P.G. performed model validation using DFT and composed the corresponding analysis. C.M.M. supervised molecular chemistry analysis of the work and assisted in writing the manuscript. E.G.S. supervised the work, provided project administration, and performed review and editing of the manuscript. All authors reviewed and approved the final version of the manuscript.
Competing interests
The Authors declare no competing interests.
{acknowledgement}
Please use “The authors thank …” rather than “The authors would like to thank …”.
The author thanks Mats Dahlgren for version one of achemso, and Donald Arseneau for the code taken from cite to move citations after punctuation. Many users have provided feedback on the class, which is reflected in all of the different demonstrations shown in this document.
{suppinfo}
The following files are available free of charge.
  • Supporting Information - FragNet: A Graph Neural Network for Molecular Property Prediction with Four Levels of Interpretability